OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). a.)-9x - 2 b.)-9x + 5 c.)-9x + 25 d.)-9x + 27

OpenStudy (anonymous):

@eliassaab can u plz help me with this one too

OpenStudy (anonymous):

f(g(x))= 9 g(x) -2= 9(-x+3) -2 Can you finish it?

OpenStudy (anonymous):

f(g(x))= 9 g(x) -2= 9(-x+3) -2=-9x + 27 -2 =9x +25